OpenStudy (anonymous):

how to solve tan(x+π)+2sin(x+π)=0

OpenStudy (akashdeepdeb):

x+pi for tan translates to tan of x x+pi for sine translates to sine of -x. so tan(x+π)+2sin(x+π)=0 = tan x - 2 sin x = 0 tan x = sinx/cos x sin x - 2sin x cos x = 0 sin x (1-2cos x) = 0 sin x =0, 1-2cos x =0 Solve them for the general solutions! Got it? :)

OpenStudy (anonymous):

yes thanks

OpenStudy (haseeb96):

since tan(x+y)=tanx+tany\1-tanxtany and sin(a+b)=sinacosb+cosasinb \[\frac{ tanx+\tan \pi }{ 1-tanxtan \pi }+2[sinxcos \pi+cosxsin \pi]\]=0 \[2\pi=360 so \pi=180\] tan180=0 , sin180=0 but cos180=-1 now put it in the given equation \[\frac{ tanx +0 }{ 1- tanx(0) }+2[sinx(-1)+cosx(0)]\]=0 \[tanx -2sinx =0\] \[\frac{ sinx }{ cosx }-2sinx=0\] sinx(1\cosx -2)=0 sinx(secx-2)=0 since,1\cosx=secx either, sinx=0 or secx+2=0 x=sin^-1(0) or secx=-2 x=0 or x=sec^-1(2) x=0 or x= 120 x=0 or x=2pi\3 this is the correct answer